OpenStudy (anonymous):

simplify: sqrt6(9-sqrt6)+6sqrt5

OpenStudy (anonymous):

funfact: \[\sqrt{ab} = \sqrt{a} \times \sqrt{b}\]

OpenStudy (jhannybean):

\[\large \sqrt{6}(9-\sqrt{6})+6 \sqrt{5}\]\[\large(9)(\sqrt{6})-(\sqrt{6})(\sqrt{6})+6\sqrt{5}\] simplify